OpenStudy (anonymous):

box has a height of 15 inches and a square base of side length 12 inches. If the larger box has a height of 20 inches, what is its volume?

OpenStudy (anonymous):

3600

OpenStudy (anonymous):

what formula did you use? i get confused when it comes to finding volumes because i know l*w= area and l*w*h= volume

OpenStudy (anonymous):

i used Volume= length*breadth*height volume=12*15*20 =3600

OpenStudy (anonymous):

i get the same too but its not part of the answer choices

OpenStudy (anonymous):

3840 cubic inches 6400 cubic inches 1620 cubic inches 5120 cubic inches

OpenStudy (anonymous):

the question is wrong or incomplete

OpenStudy (anonymous):

the question is wrong or incomplete

OpenStudy (anonymous):

the question is wrong or incomplete

OpenStudy (anonymous):

the question is wrong or incomplete

OpenStudy (anonymous):

the question is wrong or incomplete

OpenStudy (anonymous):

the question is wrong or incomplete

OpenStudy (anonymous):

the question is wrong or incomplete

OpenStudy (anonymous):

the question is wrong or incomplete

OpenStudy (anonymous):

the question is wrong or incomplete

OpenStudy (anonymous):

the question is wrong or incomplete